### v2

@@ 2 The $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol and its security under unentangled attackers
$p_{ans}=\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}]+\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}]=\frac{1}{2}\sum_{x,a\in\{0,1\}}\left\langle\psi\right|A_{a}^{x}B_{a}^{x}\left|\psi\right\rangle,$ (15)
@@
where we used the following simplified notation: when clear from the context, tensor products, identities and $\psi$ will be omitted, e.g., $\left\langle\psi\right|V_{0}^{x}\otimes A_{1}^{x}\otimes B_{1}^{x}\left|\psi\right\rangle=\langle V_{0}^{x}A_{1}^{x}B_{1}^{x}\rangle$.
This probability of answer needs to restrict the probability that the attackers are wrong. The prover’s error $p_{err}$ is such that
@@
Notice that if $p_{ans}=1$, the attackers can always attack the protocol without being caught. Using (15), the security of the protocol can be regarded as the maximum probability that the attackers can respond without being caught, and the protocol will be proven to be secure if the attackers cannot reproduce $p_{ans}\geq\eta$ for a given $p_{err}$, we formalize this idea in the following definition.
@@
We define the security region $SR$ of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  protocol as the set of pairs $(p_{err},p_{ans})\in[0,1]\times[0,1]$ for which no strategy $\mathbf{S}^{\eta}_{MoE}$ (and thus no NE strategy) exists that breaks the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  protocol with the corresponding error and response rate. A subset of $SR$ will be denoted by $SSR$. We define the attackable region $AR$ as the complementary set of the $SR$. A subset of $AR$ will be denoted as $SAR$.
Therefore, our interest relies on maximizing expression (15) over all the strategies $\mathbf{S}^{\eta}_{MoE}$ to break the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  protocol. However, unlike the set of probabilities achievable by classical physics, the set of probabilities attainable by quantum mechanics, $\mathcal{Q}$, has uncountably many extremal points, see e.g. [BCP+14], and therefore it makes the optimization problem a tough task. On the positive side, in [NPA08], Navascués, Pironio and Acín (NPA) introduced a recursive way to construct subsets ${\mathcal{Q}_{\ell}\supset\mathcal{Q}_{\ell+1}\supset\mathcal{Q}}$ for all $\ell\in\mathbb{N}$ with the property that each of them can be tested using SDP and are such that $\cap_{\ell\in\mathbb{N}}\mathcal{Q}_{\ell}=\mathcal{Q}$.
For all $a,b\in\{0,1,\perp\}$ and all $x,x^{\prime}\in\{0,1\}$, the elements $\langle A_{a}^{x}B_{b}^{x^{\prime}}\rangle$ will appear in the maximization problem solvable via SDP, and they are bounded by linear constraints given by $\mathcal{Q}_{\ell}$, see Appendix A. In addition to these constraints, we impose the additional linear constraints derived from $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, i.e., since in the protocol the verifiers abort if they receive different messages, from (5),
@@
The proof is a particular case of the proof of Proposition 4.2.
@@
Informally, Proposition 2.7 assures that between two numerical solutions for different $p_{err}$ there are no ‘abrupt jumps’, more specifically, in Fig.  2, any solution between two plotted points cannot be grater than the point in the right.
Consider the strategy $\mathbf{S}_{MoE}^{mix}|_{p}$ given by the probabilistic mixture of playing the strategy $\mathbf{S}_{\textsc{tfkw}}$ with probability $p$ and $\mathbf{S}_{guess}$ with probability $1-p$, conditioned on answering. As long as $p_{ans}<1$, for each $p$, this mixture gives a unique pair of $(p_{err},p_{ans})$ (for $p_{ans}=1$, take the minimum $p_{err}$), and we equivalently denote $\mathbf{S}_{MoE}^{mix}|_{p}$ by the corresponding $(p_{err},p_{ans})$ as $\mathbf{S}_{MoE}^{mix}|_{(p_{err},p_{ans})}$. The values of $p_{ans}$ obtained by this strategy, see continuous line in Fig. 2, provide a region where the protocol is attackable, i.e. a $SAR$. Since the SSR obtained from the second level of the NPA hierarchy and the $SAR$ obtained from $\mathcal{S}_{EoM}^{mix}|_{p}$ are such that $SSR\cup SAR=[0,1]\times[0,1]$, up to infinitesimal precision, it means that they correspond to $SR$ and $AR$, respectively, i.e., the solutions of the SDP (19) for $\ell=2$ converge to the quantum value and are tight. This means that Fig. 2 represents a full characterization of the security of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol under photon loss with attackers that do not pre-share entanglement, and the light blue region encodes all the points $(p_{err},\eta)$ where the protocol is secure. The result is summarized as follows:
Result 2.8.
Given that the (non-entangled) attackers answer with probability $\eta$ and never respond inconsistent answers, the optimal probability that they are correct in a round of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ for $\eta\in[\frac{1}{2},1]$ is given by
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}]=\cos^{2}\big{(}\frac{\pi}{8}\big{)}\eta+\sin^{2}\big{(}\frac{\pi}{8}\big{)}(1-\eta).$ (20)
@@ 2.1 Sequential repetition of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$
In every round of the protocol, the attackers will pick a strategy $S^{i}$ that can depend on the previous rounds. A strategy $S^{i}$ will induce a probability vector $\textbf{q}^{i}=(q_{\textsc{c}}^{i},q_{\perp}^{i},q_{\textsc{w}}^{i},q_{\lightning}^{i})$, where, analogous to (6), the subscripts denote the event that the random variable $\textsc{V}_{\textsc{AB}}$ takes the values ‘Correct’, ‘No photon’, ‘Wrong’ and ‘Abort’, respectively, in the $i$th round. Recall that an attack is successful if the verifiers cannot distinguish if their data came from the distribution $\textbf{p}_{hp}...\textbf{p}_{hp}$ ($r$ times) or from $\textbf{q}^{1}...\textbf{q}^{r}$.
Consider the following relaxation where the verifiers categorize wrong and abort answers identically, i.e., they count them as incorrect (i) answers. Denote by $\textbf{a}_{i}\in\{\textsc{c},\perp,\textsc{i}\}$ whether the answer they recorded in round $i$ was correct, no photon, or incorrect. Consider the following payoff function
$T_{i}(\textbf{a}_{i})=\sin^{2}\frac{\pi}{8}\textbf{1}_{\textsc{c}}(\textbf{a}_{i})-\sin^{2}\frac{\pi}{8}\textbf{1}_{\perp}(\textbf{a}_{i})-\cos^{2}\frac{\pi}{8}\textbf{1}_{\textsc{i}}(\textbf{a}_{i}),$ (21)
for every round $i$ of the protocol. Let $\Gamma_{r}=\sum_{i=1}^{r}T_{i}(\textbf{a}_{i})$ be the total ‘score’ after $r$ rounds. For an honest prover ($hp$), for every $i$,
and therefore $\mathbb{E}[\Gamma_{r}^{hp}]=r\alpha$ for simplicity we will assume the dependence on $\eta$ and $p_{err}$ implicit. Similarly, let $\Gamma_{r}^{att}$ denote the total score that the attackers ($att$) get.
@@
Let $\eta$ and $p_{err}$ be such that $\alpha>0$. Then, any sequential strategy that attackers who do not pre-share entanglement but are allowed to do LOBQC to break $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  fulfills that $\mathbb{E}[\Gamma_{r}^{att}]\leq 0$. Moreover, the probability that the attackers obtain a total score $r\alpha$, which is the expected value for an honest prover, is exponentially small:
$\mathbb{P}[\Gamma_{r}^{att}\geq r\alpha]\leq e^{-r\alpha^{2}/2}.$ (23)
@@
Figure 3: Probability simplex $\Delta_{2}$ for probabilities taking values on $\mathfrak{X}=\{\textsc{c},\perp,\textsc{i}\}$. Black dots correspond to numerical solutions of (19) for $\ell=2$. The dark region is the set of probabilities that Theorem 3.12 excludes and the black straight line is the intersection between $\Delta_{2}$ and the plane $\gamma_{\textsc{c}}p_{\textsc{c}}-\gamma_{\perp}p\gamma_{\perp}-\gamma_{\textsc{i}}p_{\textsc{i}}=0$.
@@ 3 The $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol and its security under entangled attackers
In the previous section, we have shown security for the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol. Our argument was restricted to attackers who do not pre-share entanglement, and it is well-known [KMS11] that such a protocol is broken with a single EPR pair. Recent work by Bluhm, Christandl and Speelman [BCS22], building on work by Buhrman, Fehr, Schaffner, and Speelman [BFSS13], has shown that when adding classical information to $\mathrm{QPV}_{\mathrm{BB84}}$, the $\mathrm{QPV}_{\mathrm{BB84}}^{f}$  protocol (see below), the protocol remains secure if the attackers hold less than $n/2-5$ qubits, making it secure against entangled attackers that hold an entangled state of smaller dimension. [^5] Note that this is only a lower bound. The actual best attack known requires an amount of entanglement that is exponential in the number of classical bits – it is very much possible that the protocol is better than proven. Recall that an attack on a QPV protocol can conceptually be split up into two rounds – the first in which the attackers hold part of the input and their pre-shared entangled state, and the second round in which the attackers have to respond to the closest verifier. The key to showing their result is by considering two possible joint states held by the attackers as a result of their first-round actions – one for an input where they have to measure the input qubit in the computational basis and one where they have to measure the qubit in the Hadamard basis. Then, these two states already have to be ‘far apart’ (even though Alice’s part does not depend on Bob’s input yet, and vice-versa), which by a counting argument makes it possible to obtain a bound on the size of the pre-shared state. In the spirit of applying an analogous counting argument to show security for the lossy case, we use the results of Section 2 to prove a lemma (Lemma 3.6) that allows us to apply a similar counting argument.
@@
(A round of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol). Let $n\in\mathbb{N}$, and consider a $2n$-bit boolean function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1\}$. We define one round of the lossy-function-BB84 protocol, denoted by $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$, as follows (see Fig. 4 for steps 2. and 3.):
@@
    $V_{0}$ and $V_{1}$ secretly agree on two random bit strings $x,y\in\{0,1\}^{n}$, which give a function value $f(x,y)\in\{0,1\}$. We will use function value 0 as denoting the computational basis and value 1 for the Hadamard basis. In addition, $V_{0}$ prepares the EPR pair $\left|\Omega\right\rangle=(\left|00\right\rangle+\left|11\right\rangle)/\sqrt{2}$.
@@
    $V_{0}$ sends one qubit $Q$ of $\left|\Omega\right\rangle$ and $x$ to $P$ and $V_{1}$ sends $y$ to $P$, coordinating their times so that $Q$, $x$ and $y$ arrive at $P$ at the same time. The verifiers are required to send the classical information at the speed of light, however, the quantum information can be sent arbitrarily slow  [^6] This is due to the fact that, for practical implementations, photons in an optical fiber are transmitted at a speed significantly lower than the speed of light in vacuum..
@@
    Immediately, $P$ measures $Q$ in the basis $f(x,y)$ and broadcasts her outcome to $V_{0}$ and $V_{1}$. If the photon is lost, she sends $\perp$.
@@
      both $a$ and $b$ arrive on time and $a=b$, and if
      • they are correct, the verifiers output ‘Correct’,
      • they are wrong, the verifiers output ‘Wrong’,
      • they are $\perp$, the verifiers output ‘No photon’,
@@
      if either $a$ or $b$ do not arrive on time or $a\neq b$, the verifiers output ‘Abort’.
In the same way as for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, after $r$ rounds, the verifiers accept the prover’s location if they reproduce (1). Notice that $\mathrm{QPV}_{\mathrm{BB84}}^{f}$  corresponds to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$  for $\eta=1$ and $p_{err}$=0.

### v3

@@ 2.1 Exact loss-tolerance of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$
$\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\right]=\frac{1}{2}\sum_{a,x}\mathrm{Tr}\left[\left|\psi\rangle\langle\psi\right|V^{x}_{a}\otimes A^{x}_{a}\otimes B^{x}_{a}\right]=\cos^{2}\frac{\pi}{8}.$ (19)
This strategy also gives
$\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{w}\right]=\sin^{2}\frac{\pi}{8},\hskip 28.45274pt\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\lightning\right]=0.$ (20)
[…]
Notice that if in a attack the attackers actually answer, meaning that they do not respond $\perp$’, we have
$\begin{split}\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\mid\textsc{V}_{\textsc{AB}}\neq\perp\right]+\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{w}\mid\textsc{V}_{\textsc{AB}}\neq\perp\right]+\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\lightning\mid\textsc{V}_{\textsc{AB}}\neq\perp\right]=1.\end{split}$ (21)
In fact, since for QPV we impose $\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\lightning\right]=0$, the above expression reduces to
$\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\mid\textsc{V}_{\textsc{AB}}\neq\perp\right]+\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{w}\mid\textsc{V}_{\textsc{AB}}\neq\perp\right]=1.$ (22)
We define the probability of winning, $p_{win}$ as the maximum probability of being correct conditioned on answering, i.e.  $p_{win}:=\max\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\mid\textsc{V}_{\textsc{AB}}\neq\perp\right]$, which has the interpretation of the normalized (over the conclusive answers) optimal probability of answering ‘Correct’. Showing that $p_{win}$ has a constant gap below one would imply that, over the conclusive rounds, the attackers cannot be correct as many times as they want, and if $p_{win}$ is below the expected value of an honest prover, i.e.  $\eta(1-p_{err})$, the attackers will not be able to mimic her behavior. On the other hand, for our security approach, we will consider the probability that the attackers can actually play (answer)—not answering $\perp$’—in the $\mathsf{G}^{BB84}_{\eta}$ game, i.e. given a strategy $\mathbf{S}_{\eta}=\{\left|\psi\right\rangle,A^{z}_{a},B^{z}_{b}\}_{z,a,b}$, the probability that they answer, $p_{ans}$,
$p_{ans}:=\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\right]+\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{w}\right]=\frac{1}{2}\sum_{a,x\in\{0,1\}}\left\langle\psi\right|A_{a}^{x}B_{a}^{x}\left|\psi\right\rangle,$ (23)
@@
where we used the following simplified notation: when clear from the context, tensor products, identities and $\psi$ will be omitted, e.g. $\left\langle\psi\right|V_{0}^{x}\otimes A_{1}^{x}\otimes B_{1}^{x}\left|\psi\right\rangle=\langle V_{0}^{x}A_{1}^{x}B_{1}^{x}\rangle$. Moreover, if attackers want to mimic an honest prover, they have to be consistent with the error $p_{err}$, that is,
@@
Notice that if $p_{ans}=1$ while satisfying (24), the attackers can always attack the protocol without being caught. Using (23), the security of the protocol can be regarded as the maximum probability that the attackers can respond without being caught, and the protocol will be proven to be secure if the attackers cannot reproduce $p_{ans}\geq\eta$ for a given $p_{err}$, we formalize this idea in the following definition.
@@
We define the security region $SR$ of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  protocol as the set of pairs $(p_{err},p_{ans})\in[0,1]\times[0,1]$ for which no strategy $\mathbf{S}^{\eta}_{MoE}$ (and thus no No-PE strategy) exists that breaks the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  protocol with the corresponding error and response rate.
A subset of $SR$ will be denoted by $SSR$. We define the attackable region $AR$ as the complementary set of the $SR$. A subset of $AR$ will be denoted as $SAR$.
Therefore, our interest relies on maximizing expression (23) over all the strategies $\mathbf{S}^{\eta}_{MoE}$ to break the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  protocol. However, unlike the set of probabilities achievable by classical physics, the set of probabilities attainable by quantum mechanics, $\mathcal{Q}$, has uncountably many extremal points, see e.g.  [BCP+14], and therefore it makes the optimization problem a tough task. On the positive side, in [NPA08], Navascués, Pironio and Acín (NPA) introduced a recursive way to construct subsets ${\mathcal{Q}_{\ell}\supset\mathcal{Q}_{\ell+1}\supset\mathcal{Q}}$ for all $\ell\in\mathbb{N}$ with the property that each of them can be tested using SDP and are such that $\cap_{\ell\in\mathbb{N}}\mathcal{Q}_{\ell}=\mathcal{Q}_{co}$, where $\mathcal{Q}_{co}\supset\mathcal{Q}$ is the set of probabilities obtained by Alice and Bob performing commuting measurements on a joint Hilbert space instead of tensor product measurements. For finite-dimensional Hilbert spaces, both sets are equivalent.
For all $a,b\in\{0,1,\perp\}$ and all $x,x^{\prime}\in\{0,1\}$, the elements $\langle A_{a}^{x}B_{b}^{x^{\prime}}\rangle$ will appear in the maximization problem solvable via SDP, and they are bounded by linear constraints given by $\mathcal{Q}_{\ell}$, see Appendix A. In addition to these constraints, we impose the additional linear constraints derived from $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, i.e. since in the protocol the verifiers abort if they receive different messages, from (10),
@@
The proof Proposition 2.8 is a particular case of the proof of Proposition 4.2.
@@
Informally, Proposition 2.9 assures that between two numerical solutions for different $p_{err}$ there are no ‘abrupt jumps’, more specifically, in Fig. 4, any solution between two plotted points cannot be grater than the point in the right.
Consider the strategy $\mathbf{S}_{MoE}^{mix}|_{p}$ given by the probabilistic mixture of playing the strategy $\mathbf{S}_{\textsc{tfkw}}$ with probability $p$ and $\mathbf{S}_{guess}$ with probability $1-p$, conditioned on answering. As long as $p_{ans}<1$, for each $p$, this mixture gives a unique pair of $(p_{err},p_{ans})$ (for $p_{ans}=1$, take the minimum $p_{err}$), and we equivalently denote $\mathbf{S}_{MoE}^{mix}|_{p}$ by the corresponding $(p_{err},p_{ans})$ as $\mathbf{S}_{MoE}^{mix}|_{(p_{err},p_{ans})}$. The values of $p_{ans}$ obtained by this strategy, see continuous line in Fig. 4, provide a region where the protocol is attackable, i.e. a $SAR$.
Since the SSR obtained from the second level of the NPA hierarchy and the $SAR$ obtained from $\mathbf{S}_{EoM}^{mix}|_{p}$ are such that $SSR\cup SAR=[0,1]\times[0,1]$, up to infinitesimal precision, it means that they correspond to $SR$ and $AR$, respectively, i.e. the solutions of the SDP (27) for $\ell=2$ converge to the quantum value and are tight. This means that Fig. 4 represents a full characterization of the security of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol under photon loss with attackers that do not pre-share entanglement, and the light blue region encodes all the points $(p_{err},\eta)$ where the protocol is secure. The result is summarized as follows:
Result 2.10.
In the No-PE model, if attackers answer with probability $\eta$ and never respond inconsistent answers, the optimal probability that they answer ‘Correct’ in a round of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ for $\eta\in[\frac{1}{2},1]$ is given by
$\max\Pr[\textsc{V}_{\textsc{AB}}=\textsc{c}]=\cos^{2}\big{(}\frac{\pi}{8}\big{)}\eta+\sin^{2}\big{(}\frac{\pi}{8}\big{)}(1-\eta).$ (28)
@@ 2.2 Sequential repetition of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$
Recall that in every round of the protocol, attackers will pick a strategy $\mathbf{S}^{i}_{\eta}$ that can depend on the previous rounds. Assume that the verifiers did not receive any ‘Abort’ answers, otherwise, they reject the location. Let $\Gamma_{r}^{att}$ denote the total score that the attackers ($att$) get, defined in (2). In the next theorem, we show that attackers in the No-PE model will fail the test $\mathsf{T}^{r\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$ with exponentially high probability.
@@
Consider the $r$ sequential repetition of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$. Let $\varepsilon_{\textrm{h}}>0$, $\eta$ and $p_{err}$ be such that $\alpha-\delta>0$, with $\delta=(\cos^{4}\frac{\pi}{8}\ln(1/\varepsilon_{\textrm{h}})/r)^{1/2}$. Then, any sequential strategy to attack $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  in the No-PE model fulfills that $\mathbb{E}[\Gamma_{r}^{att}]\leq 0$. Moreover, the probability that the attackers are accepted in the $\mathsf{T}^{r\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$ test is exponentially small:
$\mathrm{Pr}\left[\Gamma_{r}^{att}\geq r(\alpha-\delta)\right]\leq e^{-r(\alpha-\delta)^{2}/2}.$ (29)
Theorem 2.11 shows that there exists a test which is both complete and sound, and thus, the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ is secure for the corresponding $\eta$ and $p_{err}$. The existence of this test implies that after $r$ rounds, attackers will be caught with exponentially high probability.
The points $(\eta,p_{err})$ such that $\alpha>0$ correspond to the blue region in Fig. 4 and also below the black dots in Fig. 5—the above $\alpha-\delta$ corresponds to a small shift that can be made small increasing the number of sequential repetitions. The proof of Theorem 2.11 is a particular case of the proof of Theorem 3.13.
@@ 3 The $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol and its security under entangled attackers
In the previous section, we have shown security for the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol. Our argument was restricted to attackers who do not pre-share entanglement, and it is well-known [KMS11] that such a protocol is broken with a single EPR pair. Recent work by Bluhm, Christandl and Speelman [BCS22], building on work by Buhrman, Fehr, Schaffner, and Speelman [BFSS13], has shown that when adding classical information to $\mathrm{QPV}_{\mathrm{BB84}}$, the $\mathrm{QPV}_{\mathrm{BB84}}^{f}$  protocol (see below), the protocol remains secure if the attackers hold less than $n/2-5$ qubits, making it secure against entangled attackers that hold an entangled state of smaller dimension. [^5] Note that this is only a lower bound. The actual best attack known requires an amount of entanglement that is exponential in the number of classical bits—it is very much possible that the protocol is better than proven. Recall that an attack on a QPV protocol can conceptually be split up into two rounds—the first in which the attackers hold part of the input and their pre-shared entangled state, and the second round in which the attackers have to respond to the closest verifier. The key to showing their result is by considering two possible joint states held by the attackers as a result of their first-round actions—one for an input where they have to measure the input qubit in the computational basis and one where they have to measure the qubit in the Hadamard basis. Then, these two states already have to be ‘far apart’ (even though Alice’s part does not depend on Bob’s input yet, and vice-versa), which by a counting argument makes it possible to obtain a bound on the size of the pre-shared state. In the spirit of applying an analogous counting argument to show security for the lossy case, we use the results of Section 2 to prove a lemma (Lemma 3.6) that allows us to apply a similar counting argument.
@@
Let $n\in\mathbb{N}$, and consider a $2n$-bit boolean function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1\}$. We define one round of the lossy-function-BB84 protocol, denoted by $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$, as follows:
@@
    $V_{0}$ and $V_{1}$ secretly agree on random bit strings $x,y\in\{0,1\}^{n}$ and a bit $v\in\{0,1\}$. Then, $V_{0}$ prepares the qubit state $H^{f(x,y)}\left|v\right\rangle\in\{\left|0\right\rangle,\left|1\right\rangle,\left|+\right\rangle\,\left|-\right\rangle\}$.
@@
    $V_{0}$ sends the qubit $H^{f(x,y)}\left|v\right\rangle$, and $x$ to P, whereas $V_{1}$ sends $y$ to P, coordinating their times so that the messages arrive simultaneously to P. The classical information is required to travel at the speed of light, whereas the quantum information can be sent arbitrarily slow.
@@
    Immediately, P computes $f(x,y)$, measures the received qubit in the basis $f(x,y)$, and broadcasts her outcome, either 0 or 1, to $V_{0}$ and $V_{1}$. If she did not receive the qubit, i.e.  the photon was lost, she sends $\perp$. Therefore, the possible answers from P are $v_{\textsf{P}}\in\{0,1,\perp\}$.
@@
      $V_{0}$ and $V_{1}$ receive their respective answers at the time corresponding to the claimed location, and they are equal, i.e.  both receive the same $v_{\textsf{P}}$, then, if
      • $v_{\textsf{P}}=v$, the verifiers record ‘Correct’, denoted by ‘c’,
      • $v_{\textsf{P}}=1-v$, the verifiers record ‘Wrong’, denoted by ‘w’,
      • $v_{\textsf{P}}=\perp$, the verifiers record ‘No photon’, denoted by $\perp$’,